OpenStudy (anonymous):

"Tiffany has been dieting for 6 months. After 1 month Tiffany lost a total of 8 pounds. After 6 months, she lost a total of 25 pounds. Find the average rate of change of pounds per month." I got 3.4. Is this correct? I used the slope formula, so m=(25-8)/(6-1) m=17/5 m=3.4

Mehek (mehek14):

\(\dfrac{25-8}{6-1} = \dfrac{17}{5} = 3.4\) you are correct

OpenStudy (anonymous):

Thank you!

Mehek (mehek14):

yw \(\ddot\smile\)